How to calculate the moles of diprotic acid H2X to reach the first and second equivalence points? Then used the moles of the diprotic acid to calculate the moles in the origional 100.0 Ml of solution. remmeber that you titrated 25.0 mL of the solution.

information given:

4.9 mL of NaOH required to reach first equivalence point

9.9 mL of NaOH reuired to reach second equivalence pont

5.0421*10-4 moles of NaOH required to reach first equivalence point

1.01871*10-3 moles of NaOH required to reach second equivalence point

concentration of NaOH is .1029 M

Mass of diprotic acid used is is .2375 g
Answer #1

1) To calculate the moles of Diprotic acid required to reach the first equivalence point we have to calculate no of moles of NaOH used to reach the first equivalence point, which is 5.0421*10-4 moles whereas to reach the second equivalence point no of moles of NaOH required is 1.01871*10-3 moles. Thus the no of moles of  H2X would be half the no of moles of NaOH used

thus no of moles of  H2X used for reaching first equivalence point = 5.0421*10-4/2 = 2.521*10-4

no of moles of  H2X used for reaching second equivalence point = 1.01871*10-3/2 = 0.509*10-3

Total no of moles in 25 mL solution = 0.761*10-3

Therefore total no of moles in 100 mL solution = 0.761*10-3 *4 = 3.044*10-3